# Taxonomy and Trends in Reinforcement Learning for Robotics and Control Systems: A Structured Review

### v3

@@ Abstract
I Introduction
Reinforcement Learning (RL), a subfield of machine learning, has emerged as a powerful paradigm for solving sequential decision-making problems [1, 2]. In RL, agents learn optimal behaviors through trial-and-error interactions with dynamic environments, guided by feedback in the form of rewards or penalties [3]. This learning strategy is particularly well-suited for robotics and control systems, where agents must operate under uncertainty, interact with physical environments, and adapt to evolving tasks and conditions [4].
Traditional control methods, such as Proportional-Integral-Derivative (PID) controllers, Linear Quadratic Regulators (LQR), and Model Predictive Control (MPC), have long been the cornerstone of robotics [5]. These methods rely on accurate system models and often require manual tuning to perform well in complex, nonlinear, or changing environments [6]. Reinforcement learning, by contrast, offers the potential for model-free or partially model-based control strategies that can automatically adapt policies through direct interaction, even when the system dynamics are unknown or too complex to model precisely [7].
The confluence of RL with advances in deep learning, commonly referred to as Deep Reinforcement Learning (DRL), has enabled the application of these methods to high-dimensional and real-time control problems [8]. This has led to significant breakthroughs in robotic locomotion, manipulation, autonomous navigation, and multi-agent coordination [9]. RL algorithms have shown promise not only in simulation but increasingly in real-world robotic systems, where they can outperform hand-engineered controllers and adapt to unstructured environments [10].
Despite its growing success, the deployment of RL in physical systems presents unique challenges, such as sample inefficiency, safety constraints, sim-to-real transfer gaps, and interpretability of learned behaviors [10, 11]. Addressing these issues requires a deep integration of RL methods with domain-specific knowledge from control theory, physics-based modeling, and hardware-aware system design.
@@ I Introduction
II Background
II-A Fundamentals of Reinforcement Learning
@@
Reinforcement Learning (RL) provides a framework in which agents learn to make sequential decisions by interacting with an environment and receiving feedback in the form of rewards. This process is commonly modeled as a Markov Decision Process (MDP), defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ [11, 12], where:
@@
The goal of the agent is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative discounted return [13]:
@@
The value function $V^{\pi}(s)$ estimates the expected return from state $s$ under policy $\pi$ [14]:
@@
Similarly, the action-value function $Q^{\pi}(s,a)$ is [15]:
@@
Learning optimal policies involves estimating the optimal value functions [15]:
@@
Algorithms such as Q-learning update estimates of $Q(s,a)$ using the Bellman equation [16]:
@@
II-B Traditional Control Methods in Robotics
@@
  • PID control: Computes control inputs as a combination of proportional, integral, and derivative terms [17]:
@@
  • Linear Quadratic Regulator (LQR): Minimizes a quadratic cost function for linear systems [18]:
@@
  • Model Predictive Control (MPC): Solves an optimization problem over a finite prediction horizon using a model of the system [19]:
@@
II-C Deep Reinforcement Learning (DRL)
Deep Reinforcement Learning (DRL) integrates the decision-making framework of Reinforcement Learning with the function approximation power of deep neural networks [20]. This combination allows agents to operate in high-dimensional, continuous, or partially observed environments, making DRL particularly suited to complex robotics and control tasks such as locomotion, manipulation, and navigation [20].
@@ 1. Deep Q-Network (DQN)
The DQN algorithm approximates the Q-value function $Q(s,a;\theta)$ using a neural network with parameters $\theta$. The network is trained to minimize the Bellman error [20]:
@@ 2. Deep Deterministic Policy Gradient (DDPG)
For continuous action spaces, DDPG combines Q-learning with deterministic policy gradients. It maintains two networks [21]:
@@ 3. Twin Delayed DDPG (TD3)
TD3 improves upon DDPG by addressing the problem of overestimation bias and training instability. It introduces [22]:
@@ 4. Soft Actor-Critic (SAC)
SAC is an off-policy algorithm that maximizes a trade-off between expected return and policy entropy [23]:
@@ 5. Proximal Policy Optimization (PPO)
PPO is a policy gradient method that balances exploration and stability using a clipped surrogate objective [24]:
@@ 6. Asynchronous Advantage Actor-Critic (A3C)
A3C utilizes multiple agents in parallel environments to stabilize training. Each worker computes gradients for an actor and critic [25]:
@@ 7. Other Advanced Algorithms
  • Distributional RL (e.g., C51, QR-DQN): Models the full return distribution instead of expected value [26, 27].
  • Hierarchical RL (e.g., HIRO): Learns policies at multiple temporal scales for better abstraction and planning [28].
  • Meta-RL (e.g., MAML, PEARL): Trains agents that adapt rapidly to new tasks using limited experience [29, 30].
@@
III Taxonomy of Reinforcement Learning Applications in Robotics
Figure 2: Taxonomy of RL Applications in Robotics and Control Systems. Components (a)–(d) depict robot task categories, training structure, learning pipeline, and deployment readiness levels [31]
@@ III Taxonomy of Reinforcement Learning Applications in Robotics
The first component of the taxonomy categorizes RL applications based on robot capabilities and interaction dynamics [32]. In the domain of single-robot operation, key competencies are grouped into locomotion, navigation, stationary manipulation, and mobile manipulation. Locomotion refers to the robot’s fundamental ability to move, typically through legs, wheels, or tracks [31]. Navigation builds upon locomotion by introducing goal-directed movement in structured or unstructured environments [33]. In parallel, manipulation tasks may occur in stationary settings (e.g., robotic arms in manufacturing) or in mobile contexts, where navigation and manipulation are integrated [34].
Beyond single-robot systems, the taxonomy recognizes more complex interactive settings [31]. Human-robot interaction (HRI) encompasses tasks where humans provide feedback, demonstrations, or shared decision-making input to guide RL behavior [33]. Multi-robot interaction introduces coordination across multiple autonomous agents, requiring communication protocols, task decomposition, and policy synchronization. These two domains emphasize the growing complexity and social integration of RL-driven robotic systems [32].
@@
At the core of RL applications lies the classical agent-environment interaction cycle [34]. An RL agent receives observations from the environment and responds with an action, which leads to a new observation and a scalar reward signal [35]. This cycle allows the agent to iteratively improve its policy through trial and error, with the goal of maximizing cumulative rewards over time. The environment may be physical or simulated, and the agent’s policy can be deterministic or stochastic [36]. This formulation underpins nearly all reinforcement learning algorithms, whether value-based, policy-gradient, or actor-critic in nature [37].
@@
The third component delves into how RL agents are trained. This includes a mix of online training in simulation or physical environments, and offline learning from logged data [38]. In many scenarios, an RL agent collects experience tuples—consisting of actions, observations, rewards, and next states—which are then used to update the policy or value functions [39]. The learning process may incorporate demonstrations from expert agents or humans to improve sample efficiency and reduce unsafe exploration, especially in robotics applications with high-stakes outcomes [40]. Eventually, the result of this pipeline is a trained policy network or model that can be deployed for real-time decision making, either reactively or as part of a planning system [39].
@@
Finally, the taxonomy introduces a maturity model that categorizes RL systems based on their level of real-world deployment [38]. At level 0, policies are evaluated only in simulation, with no exposure to real-world uncertainty [39]. Level 1 introduces limited laboratory validation, while level 2 expands testing across diverse lab setups [41]. Level 3 marks the transition to real-world environments, albeit in controlled conditions [38]. Level 4 systems demonstrate robustness under more diverse and dynamic real-world scenarios [42]. Ultimately, level 5 indicates deployment in commercial or industrial products. This readiness scale is useful for evaluating how far RL-based robotic systems have progressed from conceptual prototypes to robust, deployed technologies.
IV Key Applications
@@ IV-A Locomotion
  • Bipedal and Quadrupedal Walking: Policies trained using Proximal Policy Optimization (PPO) and Soft Actor-Critic (SAC) have achieved dynamic walking and running behaviors on simulated and physical platforms such as MIT Cheetah, ANYmal, and Cassie [43].
  • Stability and Adaptability: RL policies can learn to adapt to uneven terrain, external perturbations, or joint failures without explicit reprogramming [44].
  • Hardware Deployment: Sim-to-real transfer is accomplished using domain randomization and curriculum learning to bridge the reality gap [45].
IV-B Manipulation and Grasping
@@
  • End-to-End Control: Deep RL is used to map raw images or proprioceptive signals directly to joint torques or velocities [46].
  • Contact-Rich Tasks: Algorithms like DDPG and SAC have been used to train multi-fingered robotic hands (e.g., ShadowHand) for tasks like object flipping and tool use [47].
  • Learning from Demonstrations (LfD): Combining RL with imitation learning improves sample efficiency and safety during training [34].
IV-C Autonomous Navigation
@@
  • Dynamic Environments: RL agents learn to react to moving obstacles and changing goals using visual or LiDAR-based inputs [46].
  • End-to-End Visual Navigation: CNNs trained with RL can directly output control commands from camera feeds in cluttered environments [44].
  • Multi-Modal Fusion: RL architectures fuse GPS, IMU, vision, and range sensors to learn robust navigation strategies [43].
IV-D Industrial Automation
@@
  • Robotic Assembly: RL is used for force-sensitive insertion, gear alignment, and part manipulation with real-time feedback [48].
  • Adaptive Quality Control: Inspection robots trained with RL can adapt their scanning behavior based on defect likelihood particularly with virtual reality [49].
  • Process Optimization: RL controls temperature, pressure, or conveyor belt speed to optimize energy efficiency and throughput [50].
TABLE I: Review of Reinforcement Learning Applications in Robotics and Control Systems
@@
[51] Dexterous robotic manipulation with real-world RL Human-in-the-loop RL, RLPD, vision-based policy, dual-arm tasks Industrial robot manipulation
[52] Device control through GUIs using VLMs and RL Offline + online RL, Advantage-weighted regression, AutoEval, curriculum learning Digital agents / GUI control
[53] Efficient energy management in microgrids under uncertainty Online RL, SARSA, WPESS model optimization Smart grids / energy control
[54] Robust underwater vehicle docking Adaptive Reward Shaped PPO (ARSPPO), MDP modeling, simulation + experiments Autonomous underwater vehicles (AUV)
[55] Efficient mobile robot path planning with DRL Parallel SAC (PSAC), hybrid A* path correction, embedded testing Mobile robot navigation
[56] Efficient multi-robot scheduling in smart factories Q-learning with event-triggered updates, decentralized coordination Industrial scheduling and manufacturing
[57] Safe navigation under uncertain terrain conditions Hierarchical RL, distributional RL for uncertainty modeling Autonomous ground vehicles
[58] Dynamic quadruped locomotion with learning-based torque control Actor-Critic with torque-to-command mapping, residual control Legged robotics
[59] Learning-based task planning in human-robot collaboration DRL + symbolic task representation, reward shaping Collaborative robotics
[60] Vision-based robotic bin-picking with high success rate End-to-end DRL with 3D object localization and policy fusion Industrial robotics / object manipulation
[61] Secure control of cyber–physical systems (CPS) under actuator false data injection attacks Lyapunov-based Soft Actor–Critic (SAC), proving convergence and exponential stability Cyber–physical systems, secure robot arm control, satellite attitude control
TABLE II: Review of Reinforcement Learning Applications in Robotics and Control Systems
@@
[62] Multi-agent coordination for UAV swarm inspection tasks Graph neural network-enhanced multi-agent RL (GNN-MARL) Aerial robotics, multi-UAV collaboration
[63] Adaptive control in robotic arm tasks with minimal supervision Self-supervised reinforcement learning, sparse reward optimization Industrial robot arms, adaptive manipulation
[64] Real-time control for autonomous underwater vehicles Hierarchical DDPG, safety-constrained RL, environment modeling Marine robotics, underwater exploration
[65] Energy-efficient HVAC control using meta-RL Model-free RL, meta-policy optimization Smart buildings, HVAC systems
[66] Review of RL for intelligent HVAC systems with focus on generalization Survey of model-free RL (e.g., SAC, PPO), meta-RL for better adaptability Building energy systems, HVAC control
[67] Real-world HVAC control using DRL for energy efficiency Soft Actor-Critic (SAC), model calibration with real building data Thermal building systems, DRL deployment
[31] Survey of real-world successes of DRL in robotics Comprehensive survey; analysis of policy learning, deployment techniques, and challenges Robotics (general), multi-domain DRL applications
[68] Autonomous navigation for unmanned vehicles using DRL Deep Deterministic Policy Gradient (DDPG) with Ackermann steering kinematics Autonomous driving, mobile robotics
[69] Robust quadrotor control in complex environments using DRL TD3 algorithm with environmental awareness for path tracking Aerial robotics, quadrotor control
[70] Safe and optimal energy consumption in robot-assisted logistics Safe RL using Constrained Policy Optimization (CPO) for trajectory planning Warehouse robotics, logistics automation
[71] Multi-agent collaboration for drone fleets using RL Multi-agent reinforcement learning (MARL) with reward shaping Multi-drone coordination, aerial surveillance
V Challenges in Real-World Deployment

### v4

@@ Abstract
keywords:
[…]
1 Introduction
Reinforcement Learning (RL), a subfield of machine learning, has emerged as a powerful paradigm for solving sequential decision-making problems [shakya2023reinforcement, matsuo2022deep]. In RL, agents learn optimal behaviors through trial-and-error interactions with dynamic environments, guided by feedback in the form of rewards or penalties [ladosz2022exploration]. This learning strategy is particularly well-suited for robotics and control systems, where agents must operate under uncertainty, interact with physical environments, and adapt to evolving tasks and conditions [kaufmann2024survey].
Traditional control methods, such as Proportional-Integral-Derivative (PID) controllers, Linear Quadratic Regulators (LQR), and Model Predictive Control (MPC), have long been the cornerstone of robotics [guo2025deepseek]. These methods rely on accurate system models and often require manual tuning to perform well in complex, nonlinear, or changing environments [elguea2023review]. Reinforcement learning, by contrast, offers the potential for model-free or partially model-based control strategies that can automatically adapt policies through direct interaction, even when the system dynamics are unknown or too complex to model precisely [le2022deep].
The confluence of RL with advances in deep learning, commonly referred to as Deep Reinforcement Learning (DRL), has enabled the application of these methods to high-dimensional and real-time control problems [milani2024explainable]. This has led to significant breakthroughs in robotic locomotion, manipulation, autonomous navigation, and multi-agent coordination [meyn2022control]. RL algorithms have shown promise not only in simulation but increasingly in real-world robotic systems, where they can outperform hand-engineered controllers and adapt to unstructured environments [han2023survey].
Despite its growing success, the deployment of RL in physical systems presents unique challenges, such as sample inefficiency, safety constraints, sim-to-real transfer gaps, and interpretability of learned behaviors [han2023survey, almahamid2021reinforcement]. Addressing these issues requires a deep integration of RL methods with domain-specific knowledge from control theory, physics-based modeling, and hardware-aware system design.
@@ 1 Introduction
2 Background
2.1 Fundamentals of Reinforcement Learning
@@
Reinforcement Learning (RL) provides a framework in which agents learn to make sequential decisions by interacting with an environment and receiving feedback in the form of rewards [okafor2021heuristic, okafor2021photovoltaic, okafor2021solar]. This process is commonly modeled as a Markov Decision Process (MDP), defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ [almahamid2021reinforcement, gronauer2022multi], where:
@@
The goal of the agent is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative discounted return [prudencio2023survey, adetifa2023deep]:
@@
The value function $V^{\pi}(s)$ estimates the expected return from state $s$ under policy $\pi$ [szepesvari2022algorithms, udekwe2024comparing]:
@@
Similarly, the action-value function $Q^{\pi}(s,a)$ is [laskin2021urlb, udekwe2025evaluating]:
@@
Learning optimal policies involves estimating the optimal value functions [laskin2021urlb]:
@@
Algorithms such as Q-learning update estimates of $Q(s,a)$ using the Bellman equation [adams2022survey]:
@@
2.2 Traditional Control Methods in Robotics
@@
  • PID control: Computes control inputs as a combination of proportional, integral, and derivative terms [borase2021review]:
@@
  • Linear Quadratic Regulator (LQR): Minimizes a quadratic cost function for linear systems [chacko2023lqr]:
@@
  • Model Predictive Control (MPC): Solves an optimization problem over a finite prediction horizon using a model of the system [schwenzer2021review]:
@@
2.3 Deep Reinforcement Learning (DRL)
Deep Reinforcement Learning (DRL) integrates the decision-making framework of Reinforcement Learning with the function approximation power of deep neural networks [zhao2021dynamic]. This combination allows agents to operate in high-dimensional, continuous, or partially observed environments, making DRL particularly suited to complex robotics and control tasks such as locomotion, manipulation, and navigation [zhao2021dynamic].
@@ 1. Deep Q-Network (DQN)
The DQN algorithm approximates the Q-value function $Q(s,a;\theta)$ using a neural network with parameters $\theta$. The network is trained to minimize the Bellman error [zhao2021dynamic]:
@@ 2. Deep Deterministic Policy Gradient (DDPG)
For continuous action spaces, DDPG combines Q-learning with deterministic policy gradients. It maintains two networks [bouhamed2020autonomous]:
@@ 3. Twin Delayed DDPG (TD3)
TD3 improves upon DDPG by addressing the problem of overestimation bias and training instability. It introduces [zhang2020td3]:
@@ 4. Soft Actor-Critic (SAC)
SAC is an off-policy algorithm that maximizes a trade-off between expected return and policy entropy [de2021soft]:
@@ 5. Proximal Policy Optimization (PPO)
PPO is a policy gradient method that balances exploration and stability using a clipped surrogate objective [huang2022a2c]:
@@ 6. Asynchronous Advantage Actor-Critic (A3C)
A3C utilizes multiple agents in parallel environments to stabilize training. Each worker computes gradients for an actor and critic [chen2020intelligent]:
@@ 7. Other Advanced Algorithms
  • Distributional RL (e.g., C51, QR-DQN): Models the full return distribution instead of expected value [fathan2021deep, zwetsloot2025evaluating].
  • Hierarchical RL (e.g., HIRO): Learns policies at multiple temporal scales for better abstraction and planning [zhang2021hierarchical].
  • Meta-RL (e.g., MAML, PEARL): Trains agents that adapt rapidly to new tasks using limited experience [fu2023maml2, zhu2024pearl].
@@
3 Taxonomy of Reinforcement Learning Applications in Robotics
Figure 2: Taxonomy of RL Applications in Robotics and Control Systems. Components (a)–(d) depict robot task categories, training structure, learning pipeline, and deployment readiness levels [tang2025deep]
@@ 3 Taxonomy of Reinforcement Learning Applications in Robotics
The first component of the taxonomy categorizes RL applications based on robot capabilities and interaction dynamics [garaffa2021reinforcement]. In the domain of single-robot operation, key competencies are grouped into locomotion, navigation, stationary manipulation, and mobile manipulation. Locomotion refers to the robot’s fundamental ability to move, typically through legs, wheels, or tracks [tang2025deep]. Navigation builds upon locomotion by introducing goal-directed movement in structured or unstructured environments [singh2022reinforcement]. In parallel, manipulation tasks may occur in stationary settings (e.g., robotic arms in manufacturing) or in mobile contexts, where navigation and manipulation are integrated [ibarz2021train].
Beyond single-robot systems, the taxonomy recognizes more complex interactive settings [tang2025deep]. Human-robot interaction (HRI) encompasses tasks where humans provide feedback, demonstrations, or shared decision-making input to guide RL behavior [singh2022reinforcement]. Multi-robot interaction introduces coordination across multiple autonomous agents, requiring communication protocols, task decomposition, and policy synchronization. These two domains emphasize the growing complexity and social integration of RL-driven robotic systems [garaffa2021reinforcement].
@@
At the core of RL applications lies the classical agent-environment interaction cycle [ibarz2021train]. An RL agent receives observations from the environment and responds with an action, which leads to a new observation and a scalar reward signal [korber2021comparing]. This cycle allows the agent to iteratively improve its policy through trial and error, with the goal of maximizing cumulative rewards over time. The environment may be physical or simulated, and the agent’s policy can be deterministic or stochastic [hua2021learning]. This formulation underpins nearly all reinforcement learning algorithms, whether value-based, policy-gradient, or actor-critic in nature [zhu2021deep].
@@
The third component delves into how RL agents are trained. This includes a mix of online training in simulation or physical environments, and offline learning from logged data [yang2020data]. In many scenarios, an RL agent collects experience tuples—consisting of actions, observations, rewards, and next states—which are then used to update the policy or value functions [oliff2020reinforcement]. The learning process may incorporate demonstrations from expert agents or humans to improve sample efficiency and reduce unsafe exploration, especially in robotics applications with high-stakes outcomes [wang2020mobile]. Eventually, the result of this pipeline is a trained policy network or model that can be deployed for real-time decision making, either reactively or as part of a planning system [oliff2020reinforcement].
@@
Finally, the taxonomy introduces a maturity model that categorizes RL systems based on their level of real-world deployment [yang2020data]. At level 0, policies are evaluated only in simulation, with no exposure to real-world uncertainty [oliff2020reinforcement]. Level 1 introduces limited laboratory validation, while level 2 expands testing across diverse lab setups [joshi2020robotic]. Level 3 marks the transition to real-world environments, albeit in controlled conditions [yang2020data]. Level 4 systems demonstrate robustness under more diverse and dynamic real-world scenarios [zhang2022reinforcement]. Ultimately, level 5 indicates deployment in commercial or industrial products. This readiness scale is useful for evaluating how far RL-based robotic systems have progressed from conceptual prototypes to robust, deployed technologies.
4 Key Applications
@@ 4.1 Locomotion
  • Bipedal and Quadrupedal Walking: Policies trained using Proximal Policy Optimization (PPO) and Soft Actor-Critic (SAC) have achieved dynamic walking and running behaviors on simulated and physical platforms such as MIT Cheetah, ANYmal, and Cassie [weng2021natural].
  • Stability and Adaptability: RL policies can learn to adapt to uneven terrain, external perturbations, or joint failures without explicit reprogramming [jin2020stability].
  • Hardware Deployment: Sim-to-real transfer is accomplished using domain randomization and curriculum learning to bridge the reality gap [zhao2020sim].
4.2 Manipulation and Grasping
@@
  • End-to-End Control: Deep RL is used to map raw images or proprioceptive signals directly to joint torques or velocities [zhang2021end, ajayi2025integrating].
  • Contact-Rich Tasks: Algorithms like DDPG and SAC have been used to train multi-fingered robotic hands (e.g., ShadowHand) for tasks like object flipping and tool use [spector2020deep].
  • Learning from Demonstrations (LfD): Combining RL with imitation learning improves sample efficiency and safety during training [ibarz2021train].
4.3 Autonomous Navigation
@@
  • Dynamic Environments: RL agents learn to react to moving obstacles and changing goals using visual or LiDAR-based inputs [zhang2021end, udekwe2022development].
  • End-to-End Visual Navigation: CNNs trained with RL can directly output control commands from camera feeds in cluttered environments [jin2020stability].
  • Multi-Modal Fusion: RL architectures fuse GPS, IMU, vision, and range sensors to learn robust navigation strategies [weng2021natural].
4.4 Industrial Automation
@@
  • Robotic Assembly: RL is used for force-sensitive insertion, gear alignment, and part manipulation with real-time feedback [ahn2023robotic, donatus2025multi].
  • Adaptive Quality Control: Inspection robots trained with RL can adapt their scanning behavior based on defect likelihood particularly with virtual reality [viharos2021reinforcement, udekwe2025human, udekwe2025virtual].
  • Process Optimization: RL controls temperature, pressure, or conveyor belt speed to optimize energy efficiency and throughput [petsagkourakis2020reinforcement].
Table 1: Review of Reinforcement Learning Applications in Robotics and Control Systems
@@
[luo2025precise] Dexterous robotic manipulation with real-world RL Human-in-the-loop RL, RLPD, vision-based policy, dual-arm tasks Industrial robot manipulation
[bai2024digirl] Device control through GUIs using VLMs and RL Offline + online RL, Advantage-weighted regression, AutoEval, curriculum learning Digital agents / GUI control
[meng2024online] Efficient energy management in microgrids under uncertainty Online RL, SARSA, WPESS model optimization Smart grids / energy control
[chu2025adaptive] Robust underwater vehicle docking Adaptive Reward Shaped PPO (ARSPPO), MDP modeling, simulation + experiments Autonomous underwater vehicles (AUV)
[liu2024deep] Efficient mobile robot path planning with DRL Parallel SAC (PSAC), hybrid A* path correction, embedded testing Mobile robot navigation
[xu2024reinforcement] Efficient multi-robot scheduling in smart factories Q-learning with event-triggered updates, decentralized coordination Industrial scheduling and manufacturing
[wu2024multi] Safe navigation under uncertain terrain conditions Hierarchical RL, distributional RL for uncertainty modeling Autonomous ground vehicles
[wang2024research] Dynamic quadruped locomotion with learning-based torque control Actor-Critic with torque-to-command mapping, residual control Legged robotics
[WANG2023112778] Learning-based task planning in human-robot collaboration DRL + symbolic task representation, reward shaping Collaborative robotics
[fu2023ed] Vision-based robotic bin-picking with high success rate End-to-end DRL with 3D object localization and policy fusion Industrial robotics / object manipulation
[wu2023deep] Secure control of cyber–physical systems (CPS) under actuator false data injection attacks Lyapunov-based Soft Actor–Critic (SAC), proving convergence and exponential stability Cyber–physical systems, secure robot arm control, satellite attitude control
Table 2: Review of Reinforcement Learning Applications in Robotics and Control Systems
@@
[zhou2024indoor] Multi-agent coordination for UAV swarm inspection tasks Graph neural network-enhanced multi-agent RL (GNN-MARL) Aerial robotics, multi-UAV collaboration
[zhang2024multi] Adaptive control in robotic arm tasks with minimal supervision Self-supervised reinforcement learning, sparse reward optimization Industrial robot arms, adaptive manipulation
[yang2024human] Real-time control for autonomous underwater vehicles Hierarchical DDPG, safety-constrained RL, environment modeling Marine robotics, underwater exploration
[wang2024distrl] Energy-efficient HVAC control using meta-RL Model-free RL, meta-policy optimization Smart buildings, HVAC systems
[al2024reinforcement] Review of RL for intelligent HVAC systems with focus on generalization Survey of model-free RL (e.g., SAC, PPO), meta-RL for better adaptability Building energy systems, HVAC control
[silvestri2024real] Real-world HVAC control using DRL for energy efficiency Soft Actor-Critic (SAC), model calibration with real building data Thermal building systems, DRL deployment
[tang2025deep] Survey of real-world successes of DRL in robotics Comprehensive survey; analysis of policy learning, deployment techniques, and challenges Robotics (general), multi-domain DRL applications
[xu2024autonomous] Autonomous navigation for unmanned vehicles using DRL Deep Deterministic Policy Gradient (DDPG) with Ackermann steering kinematics Autonomous driving, mobile robotics
[yin2024autonomous] Robust quadrotor control in complex environments using DRL TD3 algorithm with environmental awareness for path tracking Aerial robotics, quadrotor control
[esteso2023reinforcement] Safe and optimal energy consumption in robot-assisted logistics Safe RL using Constrained Policy Optimization (CPO) for trajectory planning Warehouse robotics, logistics automation
[zhu2023intelligent] Multi-agent collaboration for drone fleets using RL Multi-agent reinforcement learning (MARL) with reward shaping Multi-drone coordination, aerial surveillance
5 Challenges in Real-World Deployment